# UAV-enabled Computing Power Networks: Task Completion Probability Analysis

### v1

@@ I Introduction
The proliferation of latency-sensitive and computation-intensive applications, such as augmented reality and autonomous driving [Fang2025, Hu2025], has intensified the demand for ubiquitous networks that seamlessly integrate communication and distributed computing capabilities. While edge computing architectures have expanded computational capacity at the network edge, existing computing resources remain geographically fragmented, leading to resource under-utilization and creating computing power islands, known as “island effect” [Sun2024]. The International Telecommunication Union (ITU) has recently introduced the novel concept of computing power networks to address these challenges. The aim is to build a computation-driven network based on existing communication infrastructures. However, ensuring the stability and efficiency of computing power networks faces many design challenges. One major issue is the communication bottleneck in providing ubiquitous coverage with ground users (GUs) on a wide range of mobile devices and in accessing isolated computing power islands (e.g., edge servers, computing clusters, and user-provided computing nodes). Another challenge is the cost associated with the ubiquitous deployment of computing nodes (CNs) [Sun2024, ma2025uav]. How to build a cost-effective connected computing power network to reach geographically distributed computing power resources is a challenging but important design task.
@@
Existing studies on UAV-assisted multi-access edge computing/ computing power networks [Dong2024, tao2024multi, ma2025uav] typically relies on static CN accessibility models, where the number of available CNs remains fixed regardless of UAV operational positioning. This rigid assumption fundamentally constrains performance improvements when extending CNs for computing beyond geographically constrained demand areas, as it fails to use potentially abundant but spatially distributed computing resources outside the task request areas [deng2024uav]. Unlike these conventional approaches, the proposed UAV-CPN architecture exhibits two defining characteristics. First, computing power is ubiquitously available across larger geographical areas through heterogeneous CNs, ranging from vehicles (e.g., connected and autonmous vehicles or CAVs) with onboard computational capabilities [Chen2024] to dedicated CNs. Second, strategic UAV positioning (altitude and horizontal coordinates) dynamically regulates CN accessibility, a concept explored in our prior work [Deng2024], where we assume that all CNs are located within the task request area. To illustrate the altitude-dependent trade-offs, consider vertical UAV positioning. Higher altitudes improve line of sight (LoS) probability between GUs and UAVs, improving uplink reliability for GU-to-UAV task offloading and downlink reachability to remote CNs. However, increased path loss and propagation delay due to longer transmission distances may degrade signal quality and increase communication latency. This creates a non-convex optimization landscape for UAV altitude selection. For example, the altitude that maximizes uplink transmission rates for GU-to-UAV task offloading may result in suboptimal performance for the subsequent UAV-to-CN forwarding and CN computing. This is because downlink transmission rates depend on the quality of the UAV-CN communication channel, while the accessibility of CN is constrained by the UAV’s coverage area, both of which are also influenced by the UAV’s altitude. When considering horizontal UAV positioning, these interdependencies evolve into multi-dimensional spatio-temporal trade-offs. For instance, positioning closer to high-density GU areas reduces uplink delays but may limit access to high-capacity CNs beyond GU coverage. Current research inadequately addresses these compounded effects.
@@ II-B Air-ground Channel Model
Following [Mozaffari2016], we adopt a probabilistic air-ground channel model to characterize line-of-sight (LoS) and non-line-of-sight (NLoS) connectivity probabilities for both GU-to-UAV task offloading and UAV-to-CN task forwarding links.
@@ II-C Computing Model
We consider heterogeneous CNs cooperatively operated by multiple service providers, where some factors such as hardware heterogeneity, queuing delays, and I/O interference among virtual machines collectively influence computational throughput [Huang2018]. To generically model the dynamic computing capabilities of CNs, reflecting their uncertainty as well, we characterize the computation time $t_{c}$ through its cumulative distribution function (CDF):
@@ III Task completion probability analysis
In this section, we focus on a scenario where a UAV deployed at altitude $h$ above the centroid of a circular request zone with radius $R_{u}$ forwards a single GU’s computational tasks (with data size $D$) to a CN within the service zone. This single-CN assumption establishes a theoretical performance floor for the system, providing a conservative performance benchmark that intentionally excludes performance gains from multi-CN parallel processing, a common optimization strategy in existing literature [Deng2024]. To investigate the fundamental communication-computing tradeoff, we adopt idealized conditions without consideration of multi-user interference and resource contention, leaving such analyses for future extensions due to page limit.
@@ III-C Bottleneck Analysis
Based on previous analysis, we establish that task completion fails under either comm-limited or comp-limited conditions. Specifically, transmission failure occurs when CNs reside outside the communication-constrained service zone ($r_{c}\leq r_{c}^{\mathrm{max}}(r_{u})$), leading to failure in task processing during the transmission phase. For successfully transmitted tasks, the task completion probability equals the likelihood of satisfying the computing latency requirement within the residual time budget $T_{\text{res}}$. These dual constraints induce a probabilistic thinning [Lalley20XX] of the original homogeneous PPP $\Phi_{c}$, producing a thinned PPP with spatially varying density. The resultant effective CN density is formally characterized in Proposition III.1.
@@
This qualification process represents an independent thinning of the original PPP $\Phi_{c}$. By the Poisson thinning property [Lalley20XX], the resulting spatially thinned process retains the PPP property with effective density:
@@ IV Numerical results
In this section, we evaluate the task completion probability using parameters listed in Table I unless explicitly indicated. These parameters align with the scenarios where UAVs and GUs employ wireless transmission modules to support low-latency services like real-time video analytics [Fang2025]. Due to the page limit, our evaluation focuses on three critical aspects: i) the validation of theoretical models, ii) the trade-off of communication-computing resources through strategic positioning of UAV altitude, and iii) the potential for performance enhancement with more reachable CNs.
@@
Finally, we investigate the performance gain achieved through CN distribution radius expansion. Fig. 4 presents a 3D surface plot that reveals the coupled effects of CN spatial distribution and UAV altitude on task completion probability. For fixed UAV altitudes (e.g., at a UAV altitude of 300300m), extending CN distribution beyond the request zone boundaries (specifically when the distribution radius exceeds 200200m) produces substantial performance gains. Specifically, expanding the CN distribution radius from $200$ m (corresponding to our prior work [Deng2024]) to 10001000m achieves a $2.13\times$ improvement in task completion probability (46.6546.65% $\rightarrow$ 99.1499.14%), confirming our framework’s capability to leverage ubiquitous computing power distribution. For fixed CN distributions, we confirm the existence of an altitude-dependent performance maximum, demonstrating the effectiveness of dynamic CN accessibility control , which is consistent with our prior work [Deng2024]. While the UAV-CPN architecture introduces multi-dimensional coordination complexity, this proves to be essential for overcoming individual communication or computing bottlenecks, ultimately achieving superior performance.

### v2

@@ I Introduction
The proliferation of latency-sensitive and computation-intensive applications, such as augmented reality and autonomous driving [5, 6], has intensified the demand for ubiquitous networks that seamlessly integrate communication and distributed computing capabilities. While edge computing architectures have expanded computational capacity at the network edge, existing computing resources remain geographically fragmented, leading to resource under-utilization and creating computing power islands, known as “island effect” [11]. The International Telecommunication Union (ITU) has recently introduced the novel concept of computing power networks to address these challenges. The aim is to build a computation-driven network based on existing communication infrastructures. However, ensuring the stability and efficiency of computing power networks faces many design challenges. One major issue is the communication bottleneck in providing ubiquitous coverage with ground users (GUs) on a wide range of mobile devices and in accessing isolated computing power islands (e.g., edge servers, computing clusters, and user-provided computing nodes). Another challenge is the cost associated with the ubiquitous deployment of computing nodes (CNs) [11, 9]. How to build a cost-effective connected computing power network to reach geographically distributed computing power resources is a challenging but important design task.
@@
Existing studies on UAV-assisted multi-access edge computing/ computing power networks [4, 12, 9] typically relies on static CN accessibility models, where the number of available CNs remains fixed regardless of UAV operational positioning. This rigid assumption fundamentally constrains performance improvements when extending CNs for computing beyond geographically constrained demand areas, as it fails to use potentially abundant but spatially distributed computing resources outside the task request areas [3]. Unlike these conventional approaches, the proposed UAV-CPN architecture exhibits two defining characteristics. First, computing power is ubiquitously available across larger geographical areas through heterogeneous CNs, ranging from vehicles (e.g., connected and autonmous vehicles or CAVs) with onboard computational capabilities [1] to dedicated CNs. Second, strategic UAV positioning (altitude and horizontal coordinates) dynamically regulates CN accessibility, a concept explored in our prior work [2], where we assume that all CNs are located within the task request area. To illustrate the altitude-dependent trade-offs, consider vertical UAV positioning. Higher altitudes improve line of sight (LoS) probability between GUs and UAVs, improving uplink reliability for GU-to-UAV task offloading and downlink reachability to remote CNs. However, increased path loss and propagation delay due to longer transmission distances may degrade signal quality and increase communication latency. This creates a non-convex optimization landscape for UAV altitude selection. For example, the altitude that maximizes uplink transmission rates for GU-to-UAV task offloading may result in suboptimal performance for the subsequent UAV-to-CN forwarding and CN computing. This is because downlink transmission rates depend on the quality of the UAV-CN communication channel, while the accessibility of CN is constrained by the UAV’s coverage area, both of which are also influenced by the UAV’s altitude. When considering horizontal UAV positioning, these interdependencies evolve into multi-dimensional spatio-temporal trade-offs. For instance, positioning closer to high-density GU areas reduces uplink delays but may limit access to high-capacity CNs beyond GU coverage. Current research inadequately addresses these compounded effects.
@@ II-B Air-ground Channel Model
Following [10], we adopt a probabilistic air-ground channel model to characterize line-of-sight (LoS) and non-line-of-sight (NLoS) connectivity probabilities for both GU-to-UAV task offloading and UAV-to-CN task forwarding links.
@@ II-C Computing Model
We consider heterogeneous CNs cooperatively operated by multiple service providers, where some factors such as hardware heterogeneity, queuing delays, and I/O interference among virtual machines collectively influence computational throughput [7]. To generically model the dynamic computing capabilities of CNs, reflecting their uncertainty as well, we characterize the computation time $t_{c}$ through its cumulative distribution function (CDF):
@@ III Task completion probability analysis
In this section, we focus on a scenario where a UAV deployed at altitude $h$ above the centroid of a circular request zone with radius $R_{u}$ forwards a single GU’s computational tasks (with data size $D$) to a CN within the service zone. This single-CN assumption establishes a theoretical performance floor for the system, providing a conservative performance benchmark that intentionally excludes performance gains from multi-CN parallel processing, a common optimization strategy in existing literature [2]. To investigate the fundamental communication-computing tradeoff, we adopt idealized conditions without consideration of multi-user interference and resource contention, leaving such analyses for future extensions due to page limit.
@@ III-C Bottleneck Analysis
Based on previous analysis, we establish that task completion fails under either comm-limited or comp-limited conditions. Specifically, transmission failure occurs when CNs reside outside the communication-constrained service zone ($r_{c}\leq r_{c}^{\mathrm{max}}(r_{u})$), leading to failure in task processing during the transmission phase. For successfully transmitted tasks, the task completion probability equals the likelihood of satisfying the computing latency requirement within the residual time budget $T_{\text{res}}$. These dual constraints induce a probabilistic thinning [8] of the original homogeneous PPP $\Phi_{c}$, producing a thinned PPP with spatially varying density. The resultant effective CN density is formally characterized in Proposition III.1.
@@
This qualification process represents an independent thinning of the original PPP $\Phi_{c}$. By the Poisson thinning property [8], the resulting spatially thinned process retains the PPP property with effective density:
@@ IV Numerical results
In this section, we evaluate the task completion probability using parameters listed in Table I unless explicitly indicated. These parameters align with the scenarios where UAVs and GUs employ wireless transmission modules to support low-latency services like real-time video analytics [5]. Due to the page limit, our evaluation focuses on three critical aspects: i) the validation of theoretical models, ii) the trade-off of communication-computing resources through strategic positioning of UAV altitude, and iii) the potential for performance enhancement with more reachable CNs.
@@
Finally, we investigate the performance gain achieved through CN distribution radius expansion. Fig. 4 presents a 3D surface plot that reveals the coupled effects of CN spatial distribution and UAV altitude on task completion probability. For fixed UAV altitudes (e.g., at a UAV altitude of 300300m), extending CN distribution beyond the request zone boundaries (specifically when the distribution radius exceeds 200200m) produces substantial performance gains. Specifically, expanding the CN distribution radius from $200$ m (corresponding to our prior work [2]) to 10001000m achieves a $2.13\times$ improvement in task completion probability (46.6546.65% $\rightarrow$ 99.1499.14%), confirming our framework’s capability to leverage ubiquitous computing power distribution. For fixed CN distributions, we confirm the existence of an altitude-dependent performance maximum, demonstrating the effectiveness of dynamic CN accessibility control , which is consistent with our prior work [2]. While the UAV-CPN architecture introduces multi-dimensional coordination complexity, this proves to be essential for overcoming individual communication or computing bottlenecks, ultimately achieving superior performance.